Shift summary for owl 12/12/99

Jason Hamm (jhamm@physics.arizona.edu)
Sun, 12 Dec 1999 04:26:24 -0700 (MST)


Walkthrough Summary for owl shift, 12/12/99

On Walkthrough: A. Erwin, J. Hamm

02:00 AE arrives and notes that:
- cosmic run 15021 is in progress, with no audio alarms for the
hour he is present
- there's a warning from 01:48 on rib21-T temperature being
39.95, although monitoring histogram looks ok now.
- following sdaq buttons are red:
o ONSPLIT (xtape isn't running - OK)
o VOLMON CSI (ET is off - OK)
o SERVERS (ktev3 servers SEM_LIVE aren't running; turns
green later - OK)
o TRD GAS (active gas system shutdown at 01:58; Sasha L.
paged - Sasha G. will look at it on shift at 7, OK
for now)
- flow rates for all chambers lower than nominal during gas check
- everything looks ok downstairs

05:00 JCH walkthrough. Blockhouse temp OK; otherwise, nothing new
to add to Albert's report. Back to sleep...